In Jim's pocket there are nickles, dimes and quarters. There are three times as many quarters as nickles, and five more dimes than nickles. He has a total of $3.20. How many of each coin are there? Algebraic solutions only. Thank you!
Answer by richwmiller(17219)

5n+10d+25q=3.20
q=3n
d=n+5
5n+10(5+n)+25(3n)=320
n=3
q=3*3=9
d=3+5=8
d = 8, n = 3, q = 9